# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v3

@@ III.3.2 Metrics
Quantum Volume (QV): This is a metric that quantifies the largest square (equal in width and layers) quantum circuit that a quantum processor can successfully run PhysRevA.100.032328 . Each layer consists of random permutation among qubits followed by pseudo-random 2-qubit gates. After running the random quantum circuit, the measured results should be above a certain threshold to be considered successful. Consequently, higher-quality quantum gates and qubits allegedly result in a larger QV. More specifically, ${\rm QV}=2^{D}$, where $D$ is the largest square successfully run. For example, a quantum volume 256 means that a quantum processor can use a subset of 8 qubits in the processor to successfully run a circuit with 8 sequential layers. A higher QV indicates a greater capability of the quantum processor to tackle more complex tasks, demonstrating advancements in the quality of qubits and error mitigation techniques. However, one limitation of QV is that it requires knowing the probabilities of all possible outcomes, which implies a large number of measurements that scale exponentially with the number of qubits. Another limitation is that the QV protocol exclusively considers square circuits. Therefore, QV does not offer a reliable prediction of the performance of circuits that have high depth and low width, or circuits with low depth and high width. Due to the strong limitations of this metric, IBM proposed the alternative Error Per Layered Gate (EPLG) mckay2023benchmarkingquantumprocessorperformance . This metric makes use of (Clifford) randomized benchmarking to compute the layer fidelity LF, and then $\text{EPLG}=1-\text{LF}^{1/n_{2Q}}$, where $n_{2Q}$ is the number of two-qubit gates, tipically taking the value N1N-1 for a linear chain of qubits. The exponent is used to convert the aggregate process fidelity of a layer into a “per-gate” error rate by taking the geometric mean. This gives rise to the term Error Per Layered Gate.
@@
Q-Score: Q-score is an application-focused metric which measures the maximum number of variables that a quantum processor can optimize on a standard combinatorial approximate optimization problem (for instance, a Max-Cut Problem) 9459509 . Therefore, Q-Score is remarkably an application-oriented metric, since it focuses on solving a practical task. Another advantage is the versatility of the Q-score, as it can be executed and computed on any quantum processor, although in the NISQ era highly connected processors, such as the ones in trapped-ion platforms, are favored doi:10.1126/sciadv.aau0823 . On the other hand, this versatility does facilitate the comparison of devices between different computational paradigms, such as quantum annealers and gate-based vanderschoot2023qscore . Another drawback of this metric is that it does not impose any transpilation rule. This is especially important in this approximate algorithms run in NISQ processors, since the error allowed in more drastic transpilation rules can be hidden in the noise or in the approximation. Consequently, there could be strong artificial advantages of a certain processor just due to more drastic transpilation rules, so the metric is neither repeatable nor reliable.
@@
Circuit Layer Operations Per Second (CLOPS): This metric, proposed as a natural extension of Floating Point Operations per Second (FLOPS) in supercomputers, quantifies the number of QV-type layers that the quantum processor can run per second wack2021quality . This metric was proposed by IBM to measure the performance of a processor in terms of speed, complementing QV as a metric, which was focused on measuring quality. Remarkably, CLOPS include also the classical processing and control times in the calculation. Therefore, a larger CLOPS value indicates a higher speed of the quantum processor in executing layers of the same type used in the QV. Therefore, this metric suffers from limitations similar to those of QV regarding scalability, making it neither practical nor repeatable. Additionally, the metric does not naturally extend to processors that support native multi-qubit gates, such as trapped-ion hardware, as it assumes a circuit structure composed exclusively of one and two-qubit operations. Consequently, its application to architectures with fundamentally different native gate may lead to inconsistencies in performance comparisons. Finally, as CLOPS include the time corresponding to control, one can improve it by faster classical electronics, without actually improving the processor, so reliability might also be in question. Due to these limitations, IBM proposed an update of this metric called CLOPSh www.ibm.com which updates the definition of layer and now the possible two-qubit gates are no longer arbitrary, but they are only the gates which are native in the processor (i.e. between qubits which are directly connected). The reason is other gates must be decomposed into native ones, jeopardizing the definition of layer as one composed by parallel gates. Although this update is reasonable, it does not fix any of the problems associated to CLOPS.
@@
Algorithmic Qubits (AQ): Algorithmic Qubits (AQ=$N$) is an application-oriented metric that quantifies the largest quantum circuit of width $N$ and depth $N^{2}$ that a quantum processor can successfully run chen2023benchmarking . In contrast to quantum volume, the AQ metric does not concentrate on random quantum circuits but instead on six algorithmic classes, including Hamiltonian simulation, phase estimation, quantum Fourier transform, amplitude estimation, variational quantum eigensolver (VQE) simulation, and Monte Carlo sampling. The success of the quantum circuit is measured by evaluating the fidelity, which involves comparing the ideal output probability distribution to the measured output distribution on the quantum computer. It is considered successful only if it surpasses a certain threshold. Consequently, one must know a priori the outcome of the running algorithms, which is problematic for sufficiently large processors which cannot be classically simulated. Additionally, as the success is measured by the fidelity, it also shows scalability problems. As in the case of the Q-Score, no transpilation rules are imposed, which seriously affects reliability. As a positive point, this metric can be employed in different paradigms and platforms.
@@
Cross-entropy difference and cross-entropy benchmarking (XEB): This metric has been used to quantify the correspondence between the experimental and the ideal output probability distributions of a pseudo-random quantum circuit Qsupremacy , and it has been used to calculate the XEB fidelity Arute2019 . The cross-entropy difference $\Delta H(p_{\mathrm{A}})$ is defined as
@@
[…]
@@
with an error of $k/\sqrt{m}$ and $k\simeq 1$. Here, $p_{\mathrm{U}}\left(x_{j}^{\mathrm{exp}}\right)$ is the ideal probability of the $x_{j}^{\mathrm{exp}}$ bitstring computed classically. With $m\sim 10^{3}-10^{6}$ number of samples, a sufficiently accurate approximation can be achieved. The (ideal) experimental cross-entropy difference yields a value ranging from 0 to $1$, which approximates the average gate fidelity of the circuit. When the experimental output probability distribution aligns perfectly with a uniform distribution, it represents the worst-case scenario, obtaining an experimental cross entropy difference equal to zero. On the other hand, when the experimental output distribution perfectly matches the ideal distribution, the experimental cross-entropy difference is one. It is worth emphasizing that experimentally generating random circuits poses a significant challenge. Moreover, obtaining the ideal probability distribution requires classical computational resources that scale exponentially with the number of qubits, making this metric not practical. To discuss whether this metric is repeatable and reliable, it is important to distinguish between the theoretical method used to calculate $\alpha$ and the experimental implementation used to compute $\tilde{\alpha}$. In the theoretical method, the metric satisfies the repeatable and reliable characteristics. In the experimental implementation, there are two options: one can take a fixed finite set of Haar random unitaries, using always this set, or generate a new set of random unitaries every time. The first case satisfies both properties, but the second does not. However, this is more associated with the benchmarking methodology than the metric itself, so we will analyze the metric according to the theoretical method.
Hellinger distance: The Hellinger distance is a metric used to measure the similarity between two probability distributions. It has been used to quantify the capability of a quantum computer to accurately replicate the outcomes of a specified quantum circuit e24020244 . The Hellinger distance between the ideal probability distribution $P^{\mathrm{ideal}}$ and the experimental output distribution $P^{\mathrm{exp}}$ is defined by
@@
The ideal and experimental probabilities to observe the $i$-th outcome are denoted by $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$, respectively. The number of possible outcomes is given by $N=2^{n}$, where $n$ is the number of qubits. The Hellinger distance converges to 0 for identical distributions and $1$ in cases, where there is no overlap between distributions. Obtaining the ideal output distribution involves simulating the quantum circuit on a classical computer, which makes this metric not practical. The theoretical method to compute this metric ensures its repeatability and reliability, although its practical application could present challenges. Furthermore, it is important to note that the Hellinger distance, like other classical metrics such $\ell_{1}$-Norm Distance, depends on the measurement basis and therefore may fail to capture features such as phase errors.
Heavy Output Generation (HOG):
The (HOG) metric is a measure between two probability distributions, and in the context of quantum computing, it has been used to quantify the success of a pseudo-random quantum circuit executed on a quantum computer 10.5555/3135595.3135617 . If we consider a pseudo-random circuit $U$ taken from a uniform Haar distribution, the ideal and experimental output probability of each bitstring after executing $U$ is denoted by $P^{\mathrm{ideal}}$ and $P^{\mathrm{exp}}$. A bitstring $z$ is classified as heavy in $U$, if its probability $P^{\mathrm{exp}}(z)$ is greater than the median of the distribution $P^{\mathrm{ideal}}$. The probability that a sample drawn from a distribution $P^{\mathrm{exp}}$ is heavy, named as the heavy output generation probability of $P^{\mathrm{exp}}$, is given by
@@
where $N=2^{n}$, $n$ is the number of qubits, and $\delta_{i}=1$ if $i$ is heavy in $U$, and 0 otherwise. If $P^{\mathrm{exp}}$ approximates the uniform distribution, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx 1/2$; and if $P^{\mathrm{exp}}$ is close to $P^{\mathrm{ideal}}$, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx(1+\log(2))/2$. The $\mathrm{HOG}$ metric has been employed to develop a test that assesses the performance of a quantum computer as good if it exceeds $2/3$. One limitation of $\mathrm{HOG}$ is that it requires calculating the ideal probability classically, which makes this metric not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used for the cross-entropy difference and XEB metric, which means that it will satisfy both characteristics.
$\ell_{1}$-Norm Distance: The $\ell_{1}$-Norm Distance is a metric that quantifies the total difference between the output distribution of the quantum circuit produced by a quantum computer, denoted by $P^{\mathrm{exp}}$, and the probability output distribution of the ideal quantum circuit generated by a classical computer, denoted by $P^{\mathrm{exp}}$ Mills_2021 . The $\ell_{1}$- norm distance reads as
@@
where $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$ are the ideal and experimental probabilities to observe the $i$-th outcome and $N=2^{n}$ is the number of possible outcomes, with $n$ the number of qubits. The classical resources required to obtained the ideal probability distribution grow exponentially with the number of qubits, which makes this metric not practical. The theoretical method to compute it ensures its repeatability and reliability, although its practical application could present challenges.
Collision volume: This metric has been used to quantify the noise of a quantum computer using Haar random unitaries mari2023 . Collision volume involves executing a random circuit $U$, obtaining $m$ samples, and counting how many bitstrings have been sampled more than one time, what it is called collision. It can be proven that in an ideal scenario in which the quantum computer is not noisy, the number of collisions doubles the case in which the computer is completely noisy. Then, the collision volume is defined as
@@
where $m$ is the number of samples, $N=2^{n}$ with $n$ the number of qubits, and $\hat{R}$ is the observed number of collisions. The expected value of $\hat{\Delta}$ is equal to $1$ when sampling from a typical random quantum state, and 0 when sampling from a uniform distribution. Using this metric, the authors propose a test called collision volume test, which indicates that the computer has good accuracy if $\hat{\Delta}\geq 1/2$. It is important to emphasize that the collision volume metric does not require a classical computer, but it requires around $2^{n/2+5}$ samples to work, which makes it not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used in the cross-entropy difference and XEB metric, which means that it satisfy both quality attributes.
III.3.3 Benchmarks in NISQ Era

### v4

@@ 3.4.2 Metrics
Quantum Volume (QV): This is a metric that quantifies the largest square (equal in width and layers) quantum circuit that a quantum processor can successfully run [97]. Each layer consists of random permutation among qubits followed by pseudo-random 2-qubit gates. After running the random quantum circuit, the measured results should be above a certain threshold to be considered successful. Consequently, higher-quality quantum gates and qubits allegedly result in a larger QV. More specifically, ${\rm QV}=2^{D}$, where $D$ is the largest square successfully run. For example, a quantum volume 256 means that a quantum processor can use a subset of 8 qubits in the processor to successfully run a circuit with 8 sequential layers. A higher QV indicates a greater capability of the quantum processor to tackle more complex tasks, demonstrating advancements in the quality of qubits and error mitigation techniques. However, one limitation of QV is that it requires knowing the probabilities of all possible outcomes, which implies a large number of measurements that scale exponentially with the number of qubits. Another limitation is that the QV protocol exclusively considers square circuits. Therefore, QV does not offer a reliable prediction of the performance of circuits that have high depth and low width, or circuits with low depth and high width. Due to the strong limitations of this metric, IBM proposed the alternative Error Per Layered Gate (EPLG) [138]. This metric makes use of (Clifford) randomized benchmarking to compute the layer fidelity LF, and then $\text{EPLG}=1-\text{LF}^{1/n_{2Q}}$, where $n_{2Q}$ is the number of two-qubit gates, tipically taking the value N1N-1 for a linear chain of qubits. The exponent is used to convert the aggregate process fidelity of a layer into a “per-gate” error rate by taking the geometric mean. This gives rise to the term Error Per Layered Gate. As discussed above, the Quantum Volume (QV) metric is not practical, as it requires an exponential number of operations to be evaluated. It is repeatable when the same circuits are executed a sufficient number of times under comparable hardware conditions. However, it is not reliable, since it does not fully characterize the general computational performance of a quantum processor, as previously discussed. The metric is not linear, as it scales exponentially by construction. Finally, it is consistent, since the task of implementing random unitaries can be realized across different platforms.
@@
Q-Score: Q-score is an application-focused metric which measures the maximum number of variables that a quantum processor can optimize on a standard combinatorial approximate optimization problem (for instance, a Max-Cut Problem) [139]. Therefore, Q-Score is remarkably an application-oriented metric, since it focuses on solving a practical task. Another advantage is the versatility of the Q-score, as it can be executed and computed on any quantum processor, although in the NISQ era highly connected processors, such as the ones in trapped-ion platforms, are favored [140]. On the other hand, this versatility does facilitate the comparison of devices between different computational paradigms, such as quantum annealers and gate-based [141]. This metric is practical, as the comparison relies on the optimized value obtained using a heuristic method. It is also repeatable when the same compilation rules and circuits are used, provided that a sufficiently large number of measurements is performed. The metric is reliable for comparing how effectively a quantum computer optimizes specific classes of graphs using the QAOA technique; however, it is not reliable for more general graph families or for generic optimization problems addressed with this approach. The metric is not linear, since an improvement by a factor $x$ in the metric value does not imply the ability to solve the problem $x$ times better. Finally, the metric is consistent, as it does not rely on device-specific evaluation rules and can be applied across different architectures and platforms.
@@
Circuit Layer Operations Per Second (CLOPS): This metric, proposed as a natural extension of Floating Point Operations per Second (FLOPS) in supercomputers, quantifies the number of QV-type layers that the quantum processor can run per second [142]. This metric was proposed by IBM to measure the performance of a processor in terms of speed, complementing QV as a metric, which was focused on measuring quality. Remarkably, CLOPS include also the classical processing and control times in the calculation. Therefore, a larger CLOPS value indicates a higher speed of the quantum processor in executing layers of the same type used in the QV. Therefore, this metric suffers from limitations similar to those of QV regarding scalability, making it not practical. Additionally, the metric does not naturally extend to processors that support native multi-qubit gates, such as trapped-ion hardware, as it assumes a circuit structure composed exclusively of one and two-qubit operations. Consequently, its application to architectures with fundamentally different native gate may lead to inconsistencies in performance comparisons making it not consistent. Finally, as CLOPS include the time corresponding to control, one can improve it by faster classical electronics, without actually improving the processor, so reliability and linearity might also be in question. Due to these limitations, IBM proposed an update of this metric called CLOPSh [143] which updates the definition of layer and now the possible two-qubit gates are no longer arbitrary, but they are only the gates which are native in the processor (i.e. between qubits which are directly connected). The reason is other gates must be decomposed into native ones, jeopardizing the definition of layer as one composed by parallel gates. Although this update is reasonable, it does not fix any of the problems associated to CLOPS.
@@
Algorithmic Qubits (AQ): Algorithmic Qubits (AQ=$N$) is an application-oriented metric that quantifies the largest quantum circuit of width $N$ and depth $N^{2}$ that a quantum processor can successfully run [144]. In contrast to quantum volume, the AQ metric does not concentrate on random quantum circuits but instead on six algorithmic classes, including Hamiltonian simulation, phase estimation, quantum Fourier transform, amplitude estimation, variational quantum eigensolver (VQE) simulation, and Monte Carlo sampling. The success of the quantum circuit is measured by evaluating the fidelity, which involves comparing the ideal output probability distribution to the measured output distribution on the quantum computer. It is considered successful only if it surpasses a certain threshold. Consequently, one must know a priori the outcome of the running algorithms, which is problematic for sufficiently large processors which cannot be classically simulated. Additionally, as the success is measured by the fidelity, it also shows scalability problems making the metric not practical. The metric is repeatable for the same reasons as the Q-Score. However, it is not reliable for general problems, including the six instances considered here, due to limitations in achievable circuit depth. The metric is not linear, as the required circuit depth scales quadratically. Finally, the metric is consistent, since it can be applied across different computational paradigms and hardware platforms.
@@
Cross-entropy difference and cross-entropy benchmarking (XEB): This metric has been used to quantify the correspondence between the experimental and the ideal output probability distributions of a pseudo-random quantum circuit [145], and it has been used to calculate the XEB fidelity [60]. The cross-entropy difference $\Delta H(p_{\mathrm{A}})$ is defined as
@@
where $H(p_{\mathrm{1}},p_{\mathrm{2}})=-\sum^{N}_{i=1}p_{1}(x_{i})\log{p_{2}(x_{i})}$ is the cross-entropy between the probabilities distributions $p_{1}$ and $p_{2}$, and $N$ represents the total number of possible outputs. Additionally, the experimental, ideal, and uniform probability distributions are denoted by $_{\mathrm{exp}}, pp$, and $p_{\mathrm{cl}}$, respectively. In practice, we will have an experimental implementation of the Haar random unitary $U$ as $A_{\mathrm{exp(U)}}$ with an associated probability distribution $p_{\mathrm{exp}}$ and sample $S_{\mathrm{exp}}=\{x_{1}^{\mathrm{exp}},\dots x_{m}^{\mathrm{exp}}\}$. In this case, the (ideal) experimental cross-entropy difference is defined as
@@
with an error of $k/\sqrt{m}$ and $k\simeq 1$. Here, $p_{\mathrm{U}}\left(x_{j}^{\mathrm{exp}}\right)$ is the ideal probability of the $x_{j}^{\mathrm{exp}}$ bitstring computed classically. With $m\sim 10^{3}-10^{6}$ number of samples, a sufficiently accurate approximation can be achieved. The (ideal) experimental cross-entropy difference yields a value ranging from 0 to $1$, which approximates the average gate fidelity of the circuit. When the experimental output probability distribution aligns perfectly with a uniform distribution, it represents the worst-case scenario, obtaining an experimental cross entropy difference equal to zero. On the other hand, when the experimental output distribution perfectly matches the ideal distribution, the experimental cross-entropy difference is one. It is worth emphasizing that experimentally generating random circuits poses a significant challenge. Moreover, obtaining the ideal probability distribution requires classical computational resources that scale exponentially with the number of qubits, making this metric not practical. To discuss whether this metric is repeatable and reliable, it is important to distinguish between the theoretical method used to calculate $\alpha$ and the experimental implementation used to compute $\tilde{\alpha}$. In the theoretical method, the metric satisfies the repeatable and reliable characteristics. In the experimental implementation, there are two options: one can take a fixed finite set of Haar random unitaries, using always this set, or generate a new set of random unitaries every time. The first case satisfies both properties, but the second does not. However, this is more associated with the benchmarking methodology than the metric itself, so we will analyze the metric according to the theoretical method. Finally, the metric is consistent, as its definition does not depend on any specific platform or hardware architecture.
Hellinger distance: The Hellinger distance is a metric used to measure the similarity between two probability distributions. It has been used to quantify the capability of a quantum computer to accurately replicate the outcomes of a specified quantum circuit [146]. The Hellinger distance between the ideal probability distribution $P^{\mathrm{ideal}}$ and the experimental output distribution $P^{\mathrm{exp}}$ is defined by
@@
The ideal and experimental probabilities to observe the $i$-th outcome are denoted by $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$, respectively. The number of possible outcomes is given by $N=2^{n}$, where $n$ is the number of qubits. The Hellinger distance converges to 0 for identical distributions and $1$ in cases, where there is no overlap between distributions. Since computing the ideal output distribution requires classical simulation of the quantum circuit, the metric is not practical. Nevertheless, under its theoretical formulation it is repeatable and reliable, and it is consistent because its definition is independent of the underlying hardware or platform. Furthermore, it is important to note that the Hellinger distance, like other classical metrics such $\ell_{1}$-Norm Distance, depends on the measurement basis and therefore may fail to capture features such as phase errors.
Heavy Output Generation (HOG): The (HOG) metric is a measure between two probability distributions, and in the context of quantum computing, it has been used to quantify the success of a pseudo-random quantum circuit executed on a quantum computer [147]. If we consider a pseudo-random circuit $U$ taken from a uniform Haar distribution, the ideal and experimental output probability of each bitstring after executing $U$ is denoted by $P^{\mathrm{ideal}}$ and $P^{\mathrm{exp}}$. A bitstring $z$ is classified as heavy in $U$, if its probability $P^{\mathrm{exp}}(z)$ is greater than the median of the distribution $P^{\mathrm{ideal}}$. The probability that a sample drawn from a distribution $P^{\mathrm{exp}}$ is heavy, named as the heavy output generation probability of $P^{\mathrm{exp}}$, is given by
@@
where $N=2^{n}$, $n$ is the number of qubits, and $\delta_{i}=1$ if $i$ is heavy in $U$, and 0 otherwise. If $P^{\mathrm{exp}}$ approximates the uniform distribution, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx 1/2$; and if $P^{\mathrm{exp}}$ is close to $P^{\mathrm{ideal}}$, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx(1+\log(2))/2$. The $\mathrm{HOG}$ metric has been employed to develop a test that assesses the performance of a quantum computer as good if it exceeds $2/3$. One limitation of $\mathrm{HOG}$ is that it requires calculating the ideal probability classically, which makes this metric not practical. To assess whether this metric is repeatable, reliable, and consistent, we follow the same criteria used for the cross-entropy difference and XEB metrics. Under these criteria, the metric satisfies all these properties.
$\ell_{1}$-Norm Distance: The $\ell_{1}$-Norm Distance is a metric that quantifies the total difference between the output distribution of the quantum circuit produced by a quantum computer, denoted by $P^{\mathrm{exp}}$, and the probability output distribution of the ideal quantum circuit generated by a classical computer, denoted by $P^{\mathrm{exp}}$ [148]. The $\ell_{1}$- norm distance reads as
@@
where $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$ are the ideal and experimental probabilities to observe the $i$-th outcome and $N=2^{n}$ is the number of possible outcomes, with $n$ the number of qubits. The classical resources required to obtain the ideal probability distribution scale exponentially with the number of qubits, rendering this metric not practical. For the same reasons as discussed for the Hellinger distance, this metric satisfies the same set of properties.
Collision volume: This metric has been used to quantify the noise of a quantum computer using Haar random unitaries [149]. Collision volume involves executing a random circuit $U$, obtaining $m$ samples, and counting how many bitstrings have been sampled more than one time, what it is called collision. It can be proven that in an ideal scenario in which the quantum computer is not noisy, the number of collisions doubles the case in which the computer is completely noisy. Then, the collision volume is defined as
@@
where $m$ is the number of samples, $N=2^{n}$ with $n$ the number of qubits, and $\hat{R}$ is the observed number of collisions. The expected value of $\hat{\Delta}$ is equal to $1$ when sampling from a typical random quantum state, and 0 when sampling from a uniform distribution. Using this metric, the authors propose a test called collision volume test, which indicates that the computer has good accuracy if $\hat{\Delta}\geq 1/2$. It is important to emphasize that the collision volume metric does not require a classical computer, but it requires around $2^{n/2+5}$ samples to work, which makes it not practical. To determine if this metric is repeatable, reliable and consistent, we follow the same criteria used in the cross-entropy difference and XEB metric, which means that it satisfy these quality attributes.
3.4.3 Benchmarks in NISQ Era